1. The individual supply of labour is given by the following equation:
w = 8 + 0.05L
where w is the wage rate ($/hr) and L is the number of hours (per week) that a worker would choose to work.
At w = $14/hr, what is the worker’s economic rent ($/week)?
1
Expert's answer
2019-06-17T12:46:38-0400

w = 8 + 0.05L 

14 = 8+0.05L

14-8=0.05L

6=0.05L

L=120

This means that if w= $ 14 per hour, the worker is ready to work 120 hours per week, and the worker’s economic rent ($/week): 120 * 14 = 1680$